Please show work and choose correct answer
4. Assign oxidation numbers to the sulfur in each of the polyatomic ions below. Which of the following ions WILL react with acidified KMnO4 (a strong oxidizing agent) such that the purple MnO4 is reduced to almost colorless Mn+2? 2- B. C. D. E. 02- Both A and B None of the above.
Answer #1

Oxidation number of S :

A.

S2O3^2-

2X + (-2*3) =-2

2X-6=-2

2X= 4

X= +2

B.

SO4^2-

X+ (-2*4)= -2

X-8=-2

X=+6

C.

S2O7^2-

2X+(-2*7)=-2

2X -14=-2

2X= 12

X=+6

Potassium permanganate react with sulfuric acid is as follows:

4KMnO4 + 6H2SO4 ? 4MnSO4 + 2K2SO4 + 5O2 + 6H2O

Thus the correct answer is B SO4^2